# Differential Privacy for Deep Learning in Medicine

### v1

@@ 5.3 Privacy-Utility Tradeoff Results
Alsenani et al. [93] Sensor data Accuracy = $30\%$ NA $\epsilon=10$/Accuracy = $27\%$
Chin-Cheong et al. [61] EHR AUC = $80.0\%$ $\epsilon=1$/AUC = $66.1\%$ $\epsilon=10$/AUC = $67.8\%$
Daum et al. [62] Cardiac cine MRI FID = $15.4$ $\epsilon=1$/FID = $29.8$ $\epsilon=10$/FID = $26.8$
Fan et al. [66] Tabular data IBS = ~$17$ $\epsilon=1$/IBS = ~$21$ $\epsilon=8$/IBS = ~$18$
Fu et al. [94] EEG & sensor Accuracy = $99.3\%$ $\epsilon=1$/Accuracy = $98.7\%$ NA
Gong et al. [95] Time series Accuracy = $88.5\%$ $\epsilon=2$/Accuracy = $75.3\%$ $\epsilon=10$/Accuracy = $78.2\%$
Guo, Yu et al. [68] Dermoscopy Accuracy = $69.1\%$ $\epsilon=3$/Accuracy = ~$43\%$ $\epsilon=10$/Accuracy = $61.2\%$
Kaess et al. [71] CT Accuracy = $99.5\%$ NA $\epsilon=8$ /Accuracy = $97.4\%$
Mehmood et al. [74] MRI Accuracy = $88\%$ NA $\epsilon=5.6$/Accuracy = $78\%$
Ming Y et al. [96] Tabular EMR DWS =$0.0041$ $\epsilon=1$/DWS = $0.05$ $\epsilon=10$/DWS = $0.0061$
Mueller et al. [75] Image-derived graph AUC = $98.5\%$ $\epsilon=1$/AUC = $95.1\%$ $\epsilon=10$/AUC = $97.2\%$
Nguyen et al. [67] Chest X-ray Accuracy = $99\%$ $\epsilon=0.3$/Accuracy = ~$90\%$ NA
Odeyomi et al. [97] Chest X-ray Accuracy = $92\%$ $\epsilon=2$/Accuracy = $88\%$ $\epsilon=8$/Accuracy = $90\%$
Riess et al. [39] MRI Accuracy = $90.1\%$ $\epsilon=1$/Accuracy = $81.6\%$ $\epsilon=10$/Accuracy = $90\%$
Sun et al. [98] Tabular data DOP = ~$0.75$ $\epsilon=1$/DOP = ~$1.2$ $\epsilon=10$/DOP = ~$1.2$
Tang et al. [99] Tabular data F1-score = $19.6\%$ $\epsilon=1$/F1-score = $19.79\%$ NA
Tayebi Arasteh et al. [13] Chest X-ray AUC$=68-90\%$ $\epsilon\approx 1$/AUC$=67-89\%$ $\epsilon\approx 10$/AUC$=67-89\%$
Tayebi Arasteh et al. [5] Chest X-ray AUC$=89.7\%$ $\epsilon=0.5$/AUC$=84.0\%$ $\epsilon=7.9$/AUC$=87.4\%$
Tayebi Arasteh et al. [5] CT AUC$=99.7\%$ $\epsilon=0.5$/AUC$=92\%$ $\epsilon=8.0$/AUC$=99.3\%$
Tayebi Arasteh et al. [81] Pathological speech Accuracy = $99.1\%$ $\epsilon=0.9$/Accuracy = $88.3\%$ $\epsilon=7.5$/Accuracy = $95.3\%$
Wang et al. [100] Sensor data AUC = $59\%$ $\epsilon=1$/AUC = $56\%$ NA
Yan et al. [64] Chest X-ray AUC = $99\%$ $\epsilon=2$/AUC = $96.5\%$ $\epsilon=10$/AUC = $98\%$
Zhang et al. [63] DNA-SNPs Accuracy = $96\%$ $\epsilon=1$/Accuracy = ~$87\%$ NA
Zhang et al. [101] Chest X-ray Accuracy = $92.9\%$ $\epsilon=1$/Accuracy = $84.7\%$ $\epsilon=10$/Accuracy = $91.5\%$
Ziller et al. [84] Chest X-ray AUC = $96\%$ $\epsilon=0.5$/AUC = $84\%$ NA
Ziller et al. [84] Liver CT AUC = $96\%$ $\epsilon=0.5$/Dice= $94\%$ NA
@@
5.3.1 Performance at Extremely Strong Privacy Levels
At very low privacy budgets ($\epsilon\approx 1$), the magnitude of injected noise often leads to significant performance degradation. For example, Tayebi Arasteh et al. [5] observed an AUC drop from 89.7% to 84.0% on chest X-rays and from 99.7% to 92.0% on abdominal CT when applying $\epsilon=0.5$. In pathological speech classification, accuracy fell from 99.1% to 88.3% at $\epsilon=0.9$ [81], and in dermoscopy, accuracy declined from 69.1% to  43% at $\epsilon=3$ [68]. Similar trends were reported for generative tasks, such as cardiac MRI synthesis, where Fréchet inception distance (FID) worsened substantially under $\epsilon=1$ [62].
However, there are exceptions. In a multi-institutional study on domain transfer, Tayebi Arasteh et al. [13] showed that external validation performance remained nearly unchanged across $\epsilon\approx 1$–10. This suggests that DP noise may act as a regularizer, benefiting generalization even as in-domain performance drops.
5.3.2 Utility Under Strong Privacy Levels: A Modality-Specific Perspective
Most reviewed studies operate in the strong privacy regime ($\epsilon\approx 5$–10), where performance generally remains close to non-private baselines. For instance, chest X-ray AUC dropped modestly from 89.7% to 87.4% at $\epsilon=7.9$, and CT classification remained above 99% even at $\epsilon=8$ [5]. In MRI [74] and genomic data [72], performance losses at $\epsilon=10$ were similarly modest.
Nonetheless, more fragile domains show greater vulnerability to DP noise. Studies in histopathology [58], dermoscopy [68], and speech [81] report larger performance drops, reflecting higher input variance, smaller datasets, and more complex learning tasks. Tabular and sensor-based applications—e.g., in mental health or activity monitoring—also showed significant utility loss at lower $\epsilon$ values [92, 100].
These trends emphasize the importance of modality-specific tuning: modalities like chest imaging are more robust under DP-SGD, while others require additional regularization, pretraining, or architectural adaptation to maintain clinical viability.
While most studies report $\epsilon$, the value of $\delta$—which bounds the probability of privacy failure—is often missing or inconsistently stated. Best practice recommends $\delta\leq 1/n$, where $n$ is the dataset size. Many studies follow this, setting $\delta$ between $10^{-5}$ and $10^{-6}$, but others omit it or use heuristic values without justification.
Reporting gaps extend beyond privacy parameters. Performance under multiple $\epsilon$ levels is rarely benchmarked, and core implementation details—such as clipping norm, noise multiplier, accounting method (e.g., RDP or moments accountant), and number of training epochs—are frequently missing. This limits reproducibility and cross-study comparison.
This issue is particularly acute in FL settings, where DP is often applied to distributed model updates. Although FL introduces additional privacy through decentralization, few studies evaluate how DP noise affects utility across varying budgets in such settings, hindering generalizability.
To address these gaps, we recommend several reporting standards for future work. (1) Researchers should report both non-private and private model performance to provide clear context for evaluating the impact of privacy mechanisms. (2) The values of key privacy parameters, including $\epsilon$, $\delta$, and the accounting method used (e.g., RDP or moments accountant), should be explicitly stated to ensure interpretability and comparability. (3) Performance metrics should be reported across multiple $\epsilon$ levels to characterize the trade-off between privacy and utility. (4) Finally, studies should document core model and training details, including architecture type, dataset size, pretraining strategies, and other relevant hyperparameters. Adopting these practices will improve transparency, foster reproducibility, and support more rigorous benchmarking of privacy-preserving models for clinical deployment.
6 Fairness Considerations in DP-Based Medical DL
6.1 Privacy-Fairness Interactions
While DP is designed to protect individual data, its application in DL can unintentionally exacerbate performance disparities across demographic subgroups [12, 5, 102]. This is particularly concerning in healthcare, where systemic biases already exist, and privacy mechanisms that disproportionately impact underrepresented groups can further entrench inequities. Table 4 summarizes reviewed studies that assess the fairness implications of DP in medical DL, the subgroup attributes evaluated, and the metrics used.
@@ 6.1 Privacy-Fairness Interactions
Chin-Cheong et al. [61] EHR EHR generation Age, sex
Kaess et al. [71] Chest CT Classification Race, age, sex
Tayebi Arasteh et al. [13] Chest X-ray Classification Age, sex
Tayebi Arasteh et al. [5] Chest X-ray, abdominal CT Classification Age, sex
Tayebi Arasteh et al. [81] Pathological speech Neurological disorder classification Age, sex
@@
Average odds difference [103] Any Mean gap in TPR and FPR across groups Any
Disparate impact ratio[104, 105] Any Ratio of positive prediction rates Any
@@
Equal accuracy [103] Any Difference in overall accuracy across groups Any
Equal opportunity [103] Any Difference in TPR across groups Any
Equalized odds difference [103] Any Difference in both TPR and FPR Any
FDR parity [106] Any Difference in FDR across groups Any
FOR parity [106] Any Difference in FOR across groups Any
Group-wise calibration [107] Any Of predicted probabilities Any
@@
Overdiagnosis rate [108] Imaging, tabular Difference in false positive rates across groups Age, sex
Predictive equality [103] Any Difference in FPR across groups Any
PtD [109, 5] Any Positive prediction rate gap across groups Any
@@
Subgroup AUC [5] Any, especially imaging AUC reported per group to assess disparities Age, sex, race
Treatment equality [110] Any Ratio of FN to FP across groups Any
@@
Underdiagnosis rate [108] Imaging, tabular Difference in false negative rates across groups Age, sex
@@
Table 4: Summary of privacy-fairness studies and fairness evaluation metrics. The upper section lists studies that explicitly investigate fairness impacts of applying DP in medical deep learning, including use cases and demographic groups analyzed. The lower section summarizes fairness metrics reported across studies, covering various dimensions of group-level equity. ”Any” in the modalities or fairness groups column indicates that the metric is applicable across any modality or protected attribute (e.g., age, sex, race). AUC: area under the receiver operating characteristic curve; EHR: electronic health record; FDR: false discovery rate; FOR: false omission rate; FPR: false positive rate; PtD: statistical parity difference; TPR: true positive rate.
Only a handful of studies explicitly examine fairness under DP constraints [61, 71, 13, 5, 81]. These focus primarily on group-level disparities in classification tasks across attributes such as age, sex, and race. For instance, Tayebi Arasteh et al. evaluated fairness across sex and age in chest X-ray [5] and speech classification [81]; Kaess et al. [71] examined racial disparities in chest CT; and Chin-Cheong et al. [61] analyzed synthetic EHR outputs.
Findings suggest that the impact of DP is highly context-dependent. In some cases, privacy noise interacts with data imbalance to disproportionately degrade performance for certain groups—especially minorities with fewer examples. For example, DP constraints in speech classification disproportionately harmed age groups [81].
Despite these insights, most studies report fairness at only a single privacy level. Only a few works—such as [13, 5, 81]—systematically analyze fairness across multiple $\epsilon$ values, limiting our understanding of how fairness evolves as privacy constraints tighten.
Moreover, fairness is rarely a design objective. While certain configurations (e.g., ResNet backbones or group normalization) were found to reduce subgroup variance [5, 81], these choices were primarily utility-driven. Fairness-aware techniques like constraint-based learning, adversarial debiasing, or subgroup reweighting are almost entirely absent.
@@
6.2 Fairness Metrics and Their Application
Fairness auditing is essential when applying DP to medical DL, as noise can distort performance differently across groups [111]. Reviewed studies primarily assess group-level metrics based on protected attributes such as age, sex, or race. Table 4 summarizes the main metrics used.
Most fairness evaluations focus on group-level comparisons, typically assessing disparities across sex, age, or race. The most frequently reported metric is statistical parity difference (PtD) [109, 5, 13, 81, 112, 113], which measures the difference in the proportion of positive predictions between two subgroups:
@@ 6.2 Fairness Metrics and Their Application
Here, $\hat{Y}$ is the model prediction, $A$ is a protected attribute (e.g., sex), and $a$, $b$ are two group values (e.g., male and female). A related metric is the disparate impact ratio (DIR) [104, 105], which compares the rates of favorable outcomes as a ratio:
@@
Another widely used metric is the equal opportunity difference [103, 81], which focuses on differences in true positive rates (TPRs) between groups:
@@
In this case, $Y$ is the ground truth label, with $Y=1$ indicating a positive case. Equalized odds [103] extends this concept by considering both true positive and false positive rates:
$\Delta\text{EOD}=$ $\left|\Pr(\hat{Y}=1\mid Y=1,A=a)-\Pr(\hat{Y}=1\mid Y=1,A=b)\right|$
@@
Predictive equality [103], by contrast, focuses only on disparities in false positive rates:
@@
Treatment equality [110] compares the balance between false negatives (FN) and false positives (FP) across groups:
Treatment $\text{Treatment Equality}=\frac{\text{FN}_{a}/\text{FP}_{a}}{\text{FN}_{b}/\text{FP}_{b}}$ (27)
In addition to general-purpose fairness metrics, several studies employed domain-specific measures such as underdiagnosis rate [108] (differences in false negative rates across groups), overdiagnosis rate (false positives), and subgroup AUC [5, 81, 13]. While these do not require formal equations, they are particularly important in clinical settings, where different types of errors carry different consequences for patient care.
While these metrics are useful, they must be interpreted in context. For example, parity in outcomes across age groups may be inappropriate when prevalence rates differ biologically. Furthermore, DP-induced noise may distort probability calibration, reducing the reliability of threshold-based or probabilistic fairness measures.
To ensure robustness, future studies should adopt a standardized approach to fairness auditing under DP: (1) report multiple fairness metrics to capture different dimensions of disparity; (2) justify the choice of protected attributes based on clinical relevance; (3) disaggregate results by subgroup across multiple $\epsilon$ values; and (4) apply calibration or post hoc adjustments when using probability-based metrics. The development of benchmark datasets with detailed demographic annotations is also critical to advance fair, privacy-preserving AI in medicine.
7 Alternative Privacy Mechanisms Beyond DP-SGD
7.1 Federated and Mechanism-Agnostic Privacy Approaches
While DP-SGD remains the most widely used method for enforcing DP in medical DL, it is not always ideal for real-world clinical scenarios. Centralized training can be infeasible when data is decentralized, and DP-SGD often struggles with convergence or model stability under strict privacy budgets. As a result, a growing body of work explores privacy mechanisms beyond DP-SGD—including local differential privacy (LDP), exponential and Laplace mechanisms, secure aggregation protocols, and hybrid federated designs. Table 5 summarizes these approaches across various clinical tasks and data modalities.
@@ 7.1 Federated and Mechanism-Agnostic Privacy Approaches
Al Aziz et al. [92] Tabular data Classification LM, EM/LC
Almadhoun et al. [114] Genomic data Genomic privacy protection LM/NA
Alsenani et al. [93] Sensor data Detect patient activity patterns LM, GM, EM/LC
Ay et al. [115] Brain MRI Brain tumor classification NA/NA
Chen et al. [116] DNA-Seq DNA motif discovery RR/NA
Díaz et al. [117] Brain MRI AD classification GM/NA
Faisal et al. [118] Chest X-ray Pneumonia classification GM/RDP
Fan et al. [94] EEG, sensor data Freezing of gait detection LM/LC
Fang et al. [119] Tabular data synthetic data generation GM/MA
Fedeli et al. [120] Tabular data Survival prediction LM/NA
Gong et al. [95] Time series Decision support in ICUs GM/MA
Gwon et al. [96] EHR Secure data sharing RR, LM/NA
Hatamizadeh et al. [99] Tabular data Heart failure prediction NA/NA
He et al. [121] RNA-Seq Disease vs. non-disease classification GM/LC
He et al. [122] Genomic data Synthetic genomic data generation LM/NA
Honkela et al. [123] RNA-Seq Drug sensitivity prediction LM/LC
Kim et al. [124] Sensor data Health monitoring LM/LC
Kong et al. [125] Histopathology Prostate cancer diagnosis GM/NA
Li et al. [126] Pathological speech Classification (AD detection) LM, GM/NA
Li et al. [127] EHR Classification (diagnosis) LM/NA
Ming Y et al. [128] Histopathology Classification, survival prediction GM/NA
Movahedi et al. [129] Tabular data Classification LM, EM/NA
Niinimäki et al. [130] Genomic data Classification, regression LM/LC
Sun et al. [98] Tabular data Data generation GM/RDP
Tchouka et al. [131] Unstructured clinical text Multi-label ICD-10 code classification LM, EM/NA
Tramèr et al. [132] Genomic data Identifying significant SNPs LM, EM/NA
Veeraragavan et al. [133] EHR Survival Analysis LM/LC
Wang et al. [100] Sensor data Stress detection LM/NA
Wei et al. [134] Tabular data Heart disease prediction GM/LC
Wu et al. [135] Tabular data Secure data sharing RR/NA
Yuan et al. [97] Chest X-ray Pneumonia classification GM/GDP
Zhang et al. [101] Chest X-ray Pneumonia diagnosis GM/NA
Zhang et al. [136] EHR Release of medical datasets LM/LC

### v2

@@ Results
[…]
Tayebi Arasteh et al. [72arasteh2024differentialprivacyenablesfair] Pathological speech Accuracy = $99.1\%$ $\epsilon=0.9$/Accuracy = $88.3\%$ $\epsilon=7.5$/Accuracy = $95.3\%$
Wang et al. [60DifferentialPrivateFederatedTransferLearningforMentalHealthMonitoringinEverydaySettings] Sensor data AUC = $59\%$ $\epsilon=1$/AUC = $56\%$ NA
Yan et al. [24DP-SSLoRA] Chest X-ray AUC = $99\%$ $\epsilon=2$/AUC = $96.5\%$ $\epsilon=10$/AUC = $98\%$
Zhang et al. [25zhang2022secure] DNA-sequence Accuracy = $96\%$ $\epsilon=1$/Accuracy = $87\%$ NA
Zhang et al. [62BalancingCentralizedandLocalDifferentialPrivacyinPneumoniaDiagnosis] Chest X-ray Accuracy = $92.9\%$ $\epsilon=1$/Accuracy = $84.7\%$ $\epsilon=10$/Accuracy = $91.5\%$
Ziller et al. [35ziller2021medical] Chest X-ray AUC = $96\%$ $\epsilon=0.5$/AUC = $84\%$ NA
Ziller et al. [35ziller2021medical] Abdominal CT AUC = $96\%$ $\epsilon=0.5$/Dice= $94\%$ NA
@@
Fairness considerations in DP-based medical DL
While DP is designed to protect individual data, its application in DL can unintentionally exacerbate performance disparities across demographic subgroups [tran2021decision, 22tayebi2024preserving, cummings2019compatibility]. This is particularly concerning in healthcare, where systemic biases already exist, and privacy mechanisms that disproportionately impact underrepresented groups can further entrench inequities. Table 4 summarizes reviewed studies that assess the fairness implications of DP in medical DL, the subgroup attributes evaluated, and the metrics used.
@@
Chin-Cheong et al. [74chin2020generation] EHR Text/tabular data generation Age, sex
Kaess et al. [88FairandPrivateCTContrastAgentDetection] Chest CT Classification Race, age, sex
Tayebi Arasteh et al. [13TayebiDomainTransfer] Chest X-ray Classification Age, sex
Tayebi Arasteh et al. [22tayebi2024preserving] Chest X-ray, abdominal CT Classification Age, sex
Tayebi Arasteh et al. [72arasteh2024differentialprivacyenablesfair] Pathological speech Classification Age, sex
@@
Average odds difference [hardt2016equality] Any Mean gap in TPR and FPR across groups Any
Disparate impact ratio [eeoc1978uniform, feldman2015certifying] Any Ratio of PPRs Any
@@
Equal accuracy [hardt2016equality] Any Difference in overall accuracy across groups Any
Equal opportunity [hardt2016equality] Any Difference in TPR across groups Any
Equalized odds difference [hardt2016equality] Any Difference in both TPR and FPR Any
FDR parity [chouldechova2017fair] Any Difference in FDR across groups Any
FOR parity [chouldechova2017fair] Any Difference in FOR across groups Any
Group-wise calibration [pleiss2017fairness] Any Prediction-outcome gap across groups Any
@@
Overdiagnosis rate [obermeyer2019dissecting] Image, tabular Excess positive predictions across groups Age, sex
Predictive equality [hardt2016equality] Any Difference in FPR across groups Any
PtD [dwork2012fairness, 22tayebi2024preserving] Any Gap in PPRs across groups Any
@@
Subgroup AUC [22tayebi2024preserving] Any, especially image AUC reported per group to assess disparities Age, sex, race
Treatment equality [corbett2017algorithmic] Any Ratio of FN to FP across groups Any
@@
Underdiagnosis rate [obermeyer2019dissecting] Image, tabular Difference in FNRs across groups Age, sex
@@
Table 4: Summary of privacy-fairness studies and fairness evaluation metrics. The upper section lists studies that explicitly investigate fairness impacts of applying DP in medical deep learning, including use cases and demographic groups analyzed. The lower section summarizes fairness metrics reported across studies, covering various dimensions of group-level equity. ”Any” in the modalities or fairness groups column indicates that the metric is applicable across any modality or protected attribute (e.g., age, sex, race). AUC: area under the receiver operating characteristic curve; EHR: electronic health record; FDR: false discovery rate; FNR: false negative rate; FOR: false omission rate; FPR: false positive rate; PPR: positive prediction rate; PtD: statistical parity difference; TPR: true positive rate.
Only a limited number of studies explicitly examine fairness under DP constraints [74chin2020generation, 88FairandPrivateCTContrastAgentDetection, 13TayebiDomainTransfer, 22tayebi2024preserving, 72arasteh2024differentialprivacyenablesfair]. These focus primarily on group-level disparities in classification tasks across attributes such as age, sex, and race. For instance, Tayebi Arasteh et al. evaluated fairness across sex and age in chest X-ray [22tayebi2024preserving] and speech classification [72arasteh2024differentialprivacyenablesfair]; Kaess et al. [88FairandPrivateCTContrastAgentDetection] examined racial disparities in chest CT; and Chin-Cheong et al. [74chin2020generation] analyzed synthetic EHR outputs. Findings suggest that the impact of DP is highly context-dependent. In some cases, privacy noise interacts with data imbalance to disproportionately degrade performance for certain groups, especially minorities with fewer examples. For example, DP constraints in speech classification disproportionately harmed age groups [72arasteh2024differentialprivacyenablesfair]. Despite these insights, most studies report fairness at only a single privacy level. Only a few works, such as [13TayebiDomainTransfer, 22tayebi2024preserving, 72arasteh2024differentialprivacyenablesfair], systematically analyze fairness across multiple $\epsilon$ values, limiting our understanding of how fairness evolves as privacy constraints tighten. Moreover, fairness is rarely a design objective. While certain configurations (e.g., ResNet backbones or normalization with GN) were found to reduce subgroup variance [22tayebi2024preserving, 72arasteh2024differentialprivacyenablesfair], these choices were primarily utility-driven. Fairness-aware techniques like constraint-based learning, adversarial debiasing, or subgroup reweighting are almost entirely absent.
@@
Fairness auditing is essential when applying DP to medical DL, as noise can distort performance differently across groups [67ziller2024reconciling]. Reviewed studies primarily assess group-level metrics based on protected attributes such as age, sex, or race. Table 4 summarizes the main metrics used. Most fairness evaluations focus on group-level comparisons, typically assessing disparities across sex, age, or race. The most frequently reported metric is statistical parity difference (PtD) [dwork2012fairness, 22tayebi2024preserving, 13TayebiDomainTransfer, 72arasteh2024differentialprivacyenablesfair, tayebi2024addressing, 87xing2025achieving], which measures the difference in the proportion of positive predictions between two subgroups:
@@
Here, $\hat{Y}$ is the model prediction, $A$ is a protected attribute (e.g., sex), and $a$, $b$ are two group values (e.g., male and female). A related metric is the disparate impact ratio (DIR) [eeoc1978uniform, feldman2015certifying], which compares the rates of favorable outcomes as a ratio:
@@
Another widely used metric is the equal opportunity difference [hardt2016equality, 72arasteh2024differentialprivacyenablesfair], which focuses on differences in true positive rates (TPRs) between groups:
@@
In this case, $Y$ is the ground truth label, with $Y=1$ indicating a positive case. Equalized odds [hardt2016equality] extends this concept by considering both true positive and false positive rates:
$\Delta\text{EOD}=\$$\left|\Pr(\hat{Y}=1\mid Y=1,A=a)-\Pr(\hat{Y}=1\mid Y=1,A=b)\right|$
@@
Predictive equality [hardt2016equality], by contrast, focuses only on disparities in false positive rates:
@@
Treatment equality [corbett2017algorithmic] compares the balance between false negatives (FN) and false positives (FP) across groups:
$\text{Treatment Equality}=\frac{\text{FN}_{a}/\text{FP}_{a}}{\text{FN}_{b}/\text{FP}_{b}}$ (23)
In addition to general-purpose fairness metrics, several studies employed domain-specific measures such as underdiagnosis rate [obermeyer2019dissecting] (differences in false negative rates across groups), overdiagnosis rate (difference in false positive rate across groups), and subgroup AUC [22tayebi2024preserving, 72arasteh2024differentialprivacyenablesfair, 13TayebiDomainTransfer]. While these do not require formal equations, they are particularly important in clinical settings, where different types of errors carry different consequences for patient care. Other metrics such as equal accuracy [hardt2016equality], group-wise calibration [pleiss2017fairness], false discovery rate parity and false omission rate parity [chouldechova2017fair] were also used in a small subset of studies, though less frequently than PtD and equal opportunity. While these metrics are useful, they must be interpreted in context. For example, parity in outcomes across age groups may be inappropriate when prevalence rates differ biologically. Furthermore, DP-induced noise may distort probability calibration, reducing the reliability of threshold-based or probabilistic fairness measures.
To ensure robustness, future studies should adopt a standardized approach to fairness auditing under DP: (i) report multiple fairness metrics to capture different dimensions of disparity; (ii) justify the choice of protected attributes based on clinical relevance; (iii) disaggregate results by subgroup across multiple $\epsilon$ values; and (iv) apply calibration or post hoc adjustments when using probability-based metrics. The development of benchmark datasets with detailed demographic annotations is also critical to advance fair, privacy-preserving AI in medicine.
Alternative privacy mechanisms beyond DP-SGD
While DP-SGD is the dominant approach for enforcing DP in medical deep learning, several studies explored alternative mechanisms that better match real-world deployment constraints, particularly when strict privacy budgets destabilize DP-SGD training or when data cannot be centralized. These alternatives fall into four recurring patterns (summarized in Table 5).
@@
Al Aziz et al. [31AZIZ2022104113] Tabular data Classification LM, EM/LC
Almadhoun et al. [76Inferenceattacksagainstdifferentiallyprivatequery] Genomic data Attack evaluation LM/NA
Alsenani et al. [23FAItH] Sensor data Classification LM, GM, EM/LC
Ay et al. [9Dp-FedAvg] Brain MRI Classification NA/NA
Chen et al. [16DP-FLMD] DNA-sequence Pattern discovery RR/NA
Díaz et al. [85diaz2025metric] Brain MRI Classification GM/NA
Faisal et al. [52GeneratingPrivacyPreservingSyntheticMedicalData] Chest X-ray Classification GM/RDP
Fan et al. [82DP-MLD] EEG, sensor data Classification LM/LC
Fang et al. [83DP-CTGAN] Tabular data Tabular data generation GM/MA
Fedeli et al. [61fedeli2021privacy] Tabular data Regression LM/NA
Gong et al. [27gong2023federated] Time series Classification GM/MA
Gwon et al. [42GWON2024107738] EHR Data release RR, LM/NA
Hatamizadeh et al. [34tang2024personalized] Tabular data Classification NA/NA
He et al. [12DP-SL-GAN] RNA-sequence Classification GM/LC
He et al. [75Achievingdifferentialprivacyofgenomicdata] Genomic data Tabular data generation LM/NA
Honkela et al. [38honkela2018efficient] RNA-sequence Regression LM/LC
Kim et al. [77Privacy-preservingaggregationofpersonalhealthdatastreams] Sensor data Classification LM/LC
Kong et al. [4FACLCancer] Histopathology images Classification GM/NA
Li et al. [11ADDETECTOR] Pathological speech Classification LM, GM/NA
Li et al. [44LI2019103138] EHR Classification LM/NA
Ming Y et al. [40LU2022102298] Histopathology images Classification, regression GM/NA
Movahedi et al. [58EvaluatingClassifiersTrainedonDifferentiallyPrivateSyntheticHealthData] Tabular data Classification LM, EM/NA
Niinimäki et al. [68Representationtransferfordifferentiallyprivatedrugsensitivityprediction] Genomic data Classification, regression LM/LC
Sun et al. [39SUN2023104404] Tabular data Tabular data generation GM/RDP
Tchouka et al. [17NLPICD-10] Unstructured clinical text Classification LM, EM/NA
Tramèr et al. [80DifferentialPrivacywithBoundedPriors] Genomic data Feature selection LM, EM/NA
Veeraragavan et al. [81raghavan2024differentially] EHR Regression LM/LC
Wang et al. [60DifferentialPrivateFederatedTransferLearningforMentalHealthMonitoringinEverydaySettings] Sensor data Classification LM/NA
Wei et al. [29wei2024defedhdp] Tabular data Classification GM/LC
Wu et al. [79wu2021privacy] Tabular data Data release RR/NA
Yuan et al. [49yuan2019collaborative] Chest X-ray Classification GM/GDP
Zhang et al. [62BalancingCentralizedandLocalDifferentialPrivacyinPneumoniaDiagnosis] Chest X-ray Classification GM/NA
Zhang et al. [78zhang2022differential] EHR Data release LM/LC